solve: x3 +6= 2x² + 3x
Answer #1

Sour x+6=2043 x = x² - 2x - 3x + 6=0 - = x (x-2) – 3(x-2) = 0 = (1-3) (x-2)=0 ..23=0x-2 zo x=3 1-2 ==1&u=2 . = 2, 3 - 5